# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ 1 Introduction
    We develop a spatial-sign-based max-type test procedure that is robust against heavy-tailed distributions and performs well under sparse alternatives.
@@
Notations. For $k$-dimensional $\boldsymbol{a}$, we use the notation $\|\boldsymbol{a}\|$ and $\|\boldsymbol{a}\|_{\infty}$ to denote its Euclidean norm and maximum-norm respectively. We first introduce and recall some notation. For a $m\times n$ matrix $\mathbf{M}=\left(m_{j\ell}\right)_{m\times n}$, the 1- and 2-norms of $\mathbf{M}$ are $\|\mathbf{M}\|_{1}=\max_{1\leqslant\ell\leqslant n}\sum_{j=1}^{m}\left|m_{j\ell}\right|$ and $\|\mathbf{M}\|_{2}=\left\{\lambda_{\max}\left(\mathbf{M}^{\top}\mathbf{M}\right)\right\}^{1/2}$. The Frobenius norm of $\mathbf{M}$ is $\|\mathbf{M}\|_{F}=\left\{\sum_{j=1}^{m}\sum_{\ell=1}^{n}m_{j\ell}^{2}\right\}^{1/2}$. Denote $a_{n}\lesssim b_{n}$ if there exists constant $C,a_{n}\leq Cb_{n}$ and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. Let $\psi_{\alpha}(x)=\exp\left(x^{\alpha}\right)-1$ be a function defined on $[0,\infty)$ for $\alpha>0$. Then the Orlicz norm $\|\cdot\|_{\psi_{\alpha}}$ of a $\boldsymbol{X}$ is defined as $\|\boldsymbol{X}\|_{\psi_{\alpha}}=\inf\left\{t>0,\mathbb{E}\left\{\psi_{\alpha}(|\boldsymbol{X}|/t)\right\}\leqslant 1\right\}$. Let $\operatorname{tr}(\cdot)$ be a trace for matrix, $\lambda_{\text{min }}(\cdot)$ and $\lambda_{\max}(\cdot)$ be the minimum and maximum eigenvalue for symmetric martix. $\mathbf{I}_{k}$ represents a $k$-dimensional identity matrix, and $\operatorname{diag}\left\{v_{1},v_{2},\cdots,v_{k}\right\}$ represents the diagonal matrix with entries $\boldsymbol{v}=\left(v_{1},v_{2},\cdots,v_{k}\right)$. For $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$.
@@ 2 Spatial-sign based Max-type test procedure
We need the following conditions:
@@
    The $p$-dimensional vector of common factors ${\boldsymbol{f}}_{t}$ is strictly stationary and distributed independently of the errors $\varepsilon_{it^{\prime}}$, for all $i=1,\cdots,N$ and all $t,t^{\prime}=1,\cdots,T$. The number of factors $p$ is fixed and ${\boldsymbol{f}}_{t}^{\top}{\boldsymbol{f}}_{t}\leq K<+\infty$, for a constant $K$ and all $t=1,\cdots,T$. The matrix $T^{-1}({\bf 1}_{T},{\bf F})^{\top}({\bf 1}_{T},{\bf F})$ is positive definite, and as $T\to\infty$, $T^{-1}{\bf 1}_{T}^{\top}{\bf M}_{\bf F}{\bf 1}_{T}>\tau_{\min}$ for some positive constant $\tau_{\min}$.
@@
    where $\boldsymbol{W}_{t}$ is a p-dimensional random vector with independent components, $\mathbb{E}(\boldsymbol{W}_{t})=0$, $\mathbf{\Sigma}=\mathbf{\Gamma}\mathbf{\Gamma}^{\top}$, $v_{i}$ is a nonnegative univariate random variable and is independent with the spatial sign of $\boldsymbol{W}_{t}$.
@@
where $\zeta=E(r_{t}^{2})\{E(r_{t}^{-1})\}^{2}$.
We estimate $\zeta$ by $\hat{\zeta}=\widehat{E(r_{t}^{2})}\{\widehat{E(r_{t}^{-1})}\}^{2}$ and
$\widehat{E(r_{t}^{2})}=$ $T^{-1}\sum_{t=1}^{T}||\hat{{\bf U}}_{t}||^{2},$
$\widehat{E(r_{t}^{-1})}=$ $T^{-1}\sum_{t=1}^{T}||\hat{{\bf U}}_{t}||^{-1}$
Here $\hat{{\bf U}}_{t}=U\left(\hat{\bf D}^{-1/2}({\bf Z}_{t}-\hat{\boldsymbol{\theta}})\right)$.
@@
Under Condition (C1)-(C4), for some sufficiently large constant $C>0$, if $\|\boldsymbol{\alpha}\|_{\infty}\geq C\sqrt{\log N/T}$, we have the power of $T_{SM}$ goes to one.
@@ 3 Robust Cauchy combination test
    There exist $C>0$ so that $\max_{1\leq i\leq N}\sum_{j=1}^{p}\sigma_{ij}^{2}\leq$ $(\log N)^{C}$ for all $N\geq 3;N^{-1/2}(\log N)^{C}\ll\lambda_{\min}(\mathbf{R})\leq\lambda_{\max}(\mathbf{R})\ll\sqrt{N}(\log N)^{-1}$ and $\lambda_{\max}(\mathbf{R})/\lambda_{\min}(\mathbf{R})=$ $O\left(N^{\tau}\right)$ for some $\tau\in(0,1/4)$.
@@
Under Condition (C1)-(C5), we have $T_{SS}$ are asymptotically independent with $T_{SM}$ under the null hypothesis, i.e.
@@
To analysis the power performance of the new proposed CC test, we also demonstrate that $T_{SS}$ are also asymptotically independent with $T_{SM}$ under some special alternatives. Next, we consider the relationship between $T_{SS}$ and $T_{SM}$ under local alternative hypotheses:
$H_{1}:\|\boldsymbol{\alpha}\|=O\left(\zeta_{1}^{-2}NT\sigma_{SS}\right),\left\|\mathbf{R}^{1/2}\boldsymbol{\alpha}\right\|=o\left(\zeta_{1}^{-2}NT\sigma_{SS}^{2}\right)\text{ and }|\mathcal{A}|=o\left(\frac{\lambda_{\min}(\mathbf{R})\left[\operatorname{tr}\left(\mathbf{R}^{2}\right)\right]^{1/2}}{(\log N)^{C}}\right),$ (10)
@@
Under Conditions (C1)-(C5), and the alternative hypothesis (10), we have
@@ 5 Real data application
For the entire dataset ($T=165$), the $p$-values of the PY, MAX, COM, SS, SM, and CC tests are 7.70e-09, 0.0654, 1.54e-08, 9.48e-09, 0.0298, 3.60e-08, respectively. These results provide strong evidence that $\boldsymbol{\alpha}\neq\boldsymbol{0}$. Therefore, we employ a rolling window approach, applying tests to observations within each of the $165-T$ rolling windows. Each rolling window corresponds to a length-$T$ consecutive subsequence of $\{1,\ldots,165\}$, i.e., $\{\tau,\ldots,\tau+T-1\}$ for some $\tau\in\{1,\ldots,166-T\}$. Table 2 presents the rejection ratios of the tests at significance levels $\gamma=0.01,0.05$ and for three choices of $T=48,60,72$. These values of $T$ correspond to rolling windows covering 4, 5, or 6 years, respectively. Figure 3-5 show the $p$-value sequence of each test from 2005 to 2018 for $T=48,60,72$, respectively. Table 2 reveals that spatial-sign-based tests outperform corresponding least-squares-based tests, likely due to the non-normal distribution of residuals for most securities (see Figure S4 in Liu et al. (2023)). Additionally, sum-type test procedures exhibit better performance than max-type test procedures, which is unsurprising given the large number of nonzero alphas among the securities, as reported in Feng et al. (2022b). Finally, the newly proposed CC tests demonstrate nearly the best performance in most cases, highlighting the robustness of this method. These findings are consistent with theoretical results and simulation studies.
@@
Figure 3: $p$-value sequence of each test from 2005 to 2018 for US data with $T=48$.
Figure 4: $p$-value sequence of each test from 2005 to 2018 for US data with $T=60$.
Figure 5: $p$-value sequence of each test from 2005 to 2018 for US data with $T=72$.
@@ Appendix A Proofs of all theorems
Under Condition (C1)-(C4), we have $\max_{1\leq i\leq N}|\hat{d}_{i}^{2}-d_{i}^{2}|=O_{p}(\sqrt{\log N/T})$, where $\hat{{\bf D}}=\hbox{diag}\{\hat{d}_{1}^{2},\dots,\hat{d}_{N}^{2}\}$ and ${{\bf D}}=\hbox{diag}\{{d}_{1}^{2},\dots,{d}_{N}^{2}\}$.
@@
Under Condition (C1)-(C4) and the null hypothesis, when $\min(T,N)\rightarrow\infty$, $T_{SS}\mathop{\rightarrow}\limits^{d}N(0,1)$.
@@ A.1 Proof of Theorem 1
$\quad\quad-2\hat{r}^{-2}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\hat{\boldsymbol{\theta}}-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$
$\quad\quad-2\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})^{-1/2}=\mathbf{0}$
Let ${\boldsymbol{V}}_{t}={\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$. Because $W_{ij}$ satisfies the sub-exponential properties, by Theorem 2.1 in Sambale (2023), we have
$P\{\max_{1\leq t\leq T}\boldsymbol{V}_{t}^{\top}\boldsymbol{\varepsilon}{\bf D}^{-1}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}>C\mbox{tr}({\bf R})T^{-1}(1+N^{-1/3}\log N)\}$
$\leq$ $TP\{\boldsymbol{W}^{\top}(\mathbf{\Gamma}^{\top}{\bf D}^{-1}\mathbf{\Gamma}\boldsymbol{W}\otimes\boldsymbol{V}_{t}\boldsymbol{V}_{t}^{\top})\boldsymbol{W}>C\mbox{tr}({\bf R})T^{-1}(1+N^{-1/3}\log N)\}$
$\leq$ $T\exp(C^{\prime}N^{1/3}\sqrt{\log N})\rightarrow 0,$
where $\boldsymbol{W}=(W_{11},\dots,W_{1T},\dots,W_{N1},\dots,W_{NT})\in\mathcal{R}^{NT}$. By the proof of Lemma A.3 in Feng et al. (2016) and $\max_{1\leq t\leq T}\|{\bf D}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}=O_{p}(NT^{-1})$, we can obtain that $\|\hat{\boldsymbol{\theta}}\|=O_{p}(\zeta_{1}^{-1}T^{-1/2})$. According to the Taylor expansion, we can rewrite the above equation as
@@
$\quad\quad+\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}+\delta_{1t})=\mathbf{0},$
where $\delta_{1t}=O_{p}\big{\{}(-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2}+\hat{r}^{-2}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\hat{\boldsymbol{\theta}}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}+\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t})^{2}\big{\}}$.
Because ${\boldsymbol{f}}_{t}$ is independent of the errors $\varepsilon_{it}$, we have
$E(\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2})$ (1)
$=$ $E({\boldsymbol{f}}_{t}^{\top}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})$ (2)
$=$ $O(N)E({\boldsymbol{f}}_{t}^{\top}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}{\bf I}_{T}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})$ (3)
$=$ $O(NT^{-1}),$ (4)
where the last second equality holds because $\boldsymbol{\varepsilon}_{\cdot 1},\dots,\boldsymbol{\varepsilon}_{\cdot T}$ are independent and identically distributed random vectors, and $E(\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top})=E(\boldsymbol{\varepsilon}{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top})-E\{\boldsymbol{\varepsilon}({{\bf D}}^{-1}-\hat{{\bf D}}^{-1})\boldsymbol{\varepsilon}^{\top}\}=\{1+o_{p}(1)\}E(\varepsilon_{11}^{2}/d_{1}+\dots+\varepsilon_{1N}^{2}/d_{N}){\bf I}_{T}=O(N){\bf I}_{T}$. Moreover, $\hat{r}^{-2}_{t}=\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}=\|{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}(1+\|(\hat{{\bf D}}^{-1/2}-{\bf D}^{-1/2}){\bf D}^{1/2}{\bf U}_{t}\|^{2}-2{\bf U}_{t}^{\top}(\hat{{\bf D}}^{-1/2}-{\bf D}^{-1/2}){\bf D}^{1/2}{\bf U}_{t})^{-1}={\color[rgb]{0,0,0}\|{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}}\times\{1+o_{p}(1)\}=O_{p}(N^{-1})$. Hence, we have $\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}=O(T^{-1})$. Combined the above results and Condition (C3), we have $\delta_{1t}=O_{p}(T^{-1})$. Then, we have

### v2

@@ 1 Introduction
    We develop a spatial-sign-based max-type test procedure that is robust against heavy-tailed distributions and performs well under sparse alternatives. The theoretical findings are not insignificant. Contrary to the typical mean testing problem, we must address the factors that necessitate rescaling the test statistic, a task that considerably increases the complexity of the proofs.
@@
Notations. For $k$-dimensional vector $\boldsymbol{a}$, we use the notation $\|\boldsymbol{a}\|$ and $\|\boldsymbol{a}\|_{\infty}$ to denote its Euclidean norm and maximum-norm respectively. For a $m\times n$ matrix $\mathbf{M}=\left(m_{j\ell}\right)_{m\times n}$, the 1- and 2-norms of $\mathbf{M}$ are $\|\mathbf{M}\|_{1}=\max_{1\leqslant\ell\leqslant n}\sum_{j=1}^{m}\left|m_{j\ell}\right|$ and $\|\mathbf{M}\|_{2}=\left\{\lambda_{\max}\left(\mathbf{M}^{\top}\mathbf{M}\right)\right\}^{1/2}$. The Frobenius norm of $\mathbf{M}$ is $\|\mathbf{M}\|_{F}=\left\{\sum_{j=1}^{m}\sum_{\ell=1}^{n}m_{j\ell}^{2}\right\}^{1/2}$. Denote $a_{n}\lesssim b_{n}$ if there exists constant $C,a_{n}\leq Cb_{n}$ and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. For two sequences of numbers $\left\{a_{n}\geq 0;n\geq 1\right\}$ and $\left\{b_{n}>0;n\geq 1\right\}$, we write $a_{n}\ll b_{n}$ if $\lim_{n\rightarrow\infty}\frac{a_{n}}{b_{n}}=0$. The following assumption will be imposed. Let $\psi_{\alpha}(x)=\exp\left(x^{\alpha}\right)-1$ be a function defined on $[0,\infty)$ for $\alpha>0$. Then the Orlicz norm $\|\cdot\|_{\psi_{\alpha}}$ of a $\boldsymbol{X}$ is defined as $\|\boldsymbol{X}\|_{\psi_{\alpha}}=\inf\left\{t>0,\mathbb{E}\left\{\psi_{\alpha}(|\boldsymbol{X}|/t)\right\}\leqslant 1\right\}$. Let $\operatorname{tr}(\cdot)$ be a trace for matrix, $\lambda_{\text{min }}(\cdot)$ and $\lambda_{\max}(\cdot)$ be the minimum and maximum eigenvalue for symmetric martix. $\mathbf{I}_{k}$ represents a $k$-dimensional identity matrix, and $\operatorname{diag}\left\{v_{1},v_{2},\cdots,v_{k}\right\}$ represents the diagonal matrix with entries $\boldsymbol{v}=\left(v_{1},v_{2},\cdots,v_{k}\right)$. For $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$.
@@ 2 Spatial-sign based Max-type test procedure
Let $\mathcal{F}_{-\infty}^{0}$ and $\mathcal{F}_{k}^{\infty}$ denote the $\sigma$-algebras generated by $\big{\{}{\boldsymbol{f}}_{t}:-\infty\leq t\leq 0\big{\}}$ and $\big{\{}{\boldsymbol{f}}_{t}:k\leq t\leq\infty\big{\}}$, respectively. Let $\alpha(k)\equiv\sup_{A\in\mathcal{F}_{-\infty}^{0},B\in\mathcal{F}_{k}^{\infty}}|P(A)P(B)-P(AB)|$ denote the $\alpha$-mixing coefficient. We need the following conditions:
@@
    The $p$-dimensional vector of common factors ${\boldsymbol{f}}_{t}$ is strictly stationary and distributed independently of the errors $\varepsilon_{it^{\prime}}$, for all $i=1,\cdots,N$ and all $t,t^{\prime}=1,\cdots,T$. The number of factors $p$ is fixed and ${\boldsymbol{f}}_{t}^{\top}{\boldsymbol{f}}_{t}\leq K<+\infty$, for a constant $K$ and all $t=1,\cdots,T$. The matrix $T^{-1}({\bf 1}_{T},{\bf F})^{\top}({\bf 1}_{T},{\bf F})$ is positive definite, and as $T\to\infty$, $T^{-1}{\bf 1}_{T}^{\top}{\bf M}_{\bf F}{\bf 1}_{T}>\tau_{\min}$ for some positive constant $\tau_{\min}$. There exist $c_{2},c_{3}>0$, such that for any positive integer $k$, the $\alpha$-mixing coefficient $\alpha(k)\leq\exp(-c_{2}k^{c_{3}})$.
@@
    where $\boldsymbol{W}_{t}$ is a p-dimensional random vector with independent components, $\mathbb{E}(\boldsymbol{W}_{t})=0$, $\mathbf{\Sigma}=\mathbf{\Gamma}\mathbf{\Gamma}^{\top}$, $v_{t}$ is a nonnegative univariate random variable and is independent with the spatial sign of $\boldsymbol{W}_{t}$ and $0<E(v_{t}^{2})<\infty$.
@@
where $\zeta=N\{E(r_{t}^{-1})\}^{2}/\eta_{\omega}$, $\eta_{\omega}=1-2\eta E(r_{t}^{-1})E(r_{t})+\eta E(r_{s}^{-2})E(r_{t}^{2})$ and $T^{-1}\omega_{T}\mathop{\rightarrow}\limits^{p}\omega=1-\eta$.
We estimate $\zeta$ by
$\hat{\zeta}=\frac{N\{\widehat{E(r_{t}^{-1})}\}^{2}}{1-2T^{-1}{\bf 1}_{T}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}{\bf 1}_{T}\widehat{E(r_{t}^{-1})}\widehat{E(r_{t})}+T^{-1}{\bf 1}_{T}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}{\bf 1}_{T}\widehat{E(r_{t}^{2})}\widehat{E(r_{t}^{-2})}}$
and
$\widehat{E(r_{t}^{2})}=$ $T^{-1}\sum_{t=1}^{T}||\tilde{\boldsymbol{\varepsilon}}_{t\cdot}||^{2},\,\,\,$ $\widehat{E(r_{t}^{-1})}=T^{-1}\sum_{t=1}^{T}||\tilde{\boldsymbol{\varepsilon}}_{t\cdot}||^{-1},$
$\widehat{E(r_{t})}=$ $T^{-1}\sum_{t=1}^{T}||\tilde{\boldsymbol{\varepsilon}}_{t\cdot}||,\,\,\,$ $\widehat{E(r_{t}^{-2})}=T^{-1}\sum_{t=1}^{T}||\tilde{\boldsymbol{\varepsilon}}_{t\cdot}||^{-2}$
Here $\tilde{\boldsymbol{\varepsilon}}_{t\cdot}=\hat{\bf D}^{-1/2}({\bf Z}_{t}-\hat{\boldsymbol{\theta}})$.
@@
Under Conditions (C1)-(C4), for some sufficiently large constant $C>0$, if $\|\boldsymbol{\alpha}\|_{\infty}\geq C\sqrt{\log N/T}$, $\|\boldsymbol{\alpha}\|^{2}=O(NT^{-1}\log N)$ and $\lambda_{\max}(\mathbf{R})=o(N/(\log N)^{2})$, we have the power of $T_{SM}$ goes to one.
@@ 3 Robust Cauchy combination test
    $\operatorname{tr}\left(\mathbf{R}^{4}\right)=o\left\{\operatorname{tr}^{2}\left(\mathbf{R}^{2}\right)\right\}$; $T^{-2}N^{2}/\operatorname{tr}\left(\mathbf{R}^{2}\right)=O(1)$; $\operatorname{tr}\left(\mathbf{R}^{2}\right)-N=o\left(T^{-1}N^{2}\right)$.
  • (C6)
    There exist $C>0$ so that $\max_{1\leq i\leq N}\sum_{j=1}^{N}\sigma_{ij}^{2}\leq$ $(\log N)^{C}$ for all $N\geq 3;N^{-1/2}(\log N)^{C}\ll\lambda_{\min}(\mathbf{R})\leq\lambda_{\max}(\mathbf{R})\ll\sqrt{N}(\log N)^{-1}$ and $\lambda_{\max}(\mathbf{R})/\lambda_{\min}(\mathbf{R})=$ $O\left(N^{\tau}\right)$ for some $\tau\in(0,1/4)$.
Note that Condition (C5) is identical to Assumption (A3) in Liu et al. (2023), while Condition (C6) is identical to Assumption (2.3) in Feng et al. (2022a).
@@
Under Conditions (C1)-(C6) and $\log N=o({T^{1/10}})$, we have $T_{SS}$ is asymptotically independent with $T_{SM}$ under the null hypothesis, i.e.
@@
To analysis the power performance of the new proposed CC test, we also demonstrate that $T_{SS}$ are also asymptotically independent with $T_{SM}$ under some special alternatives. We consider the relationship between $T_{SS}$ and $T_{SM}$ under local alternative hypothesis:
$H_{1}:\|\boldsymbol{\alpha}\|^{2}=O\left(\zeta_{1}^{-2}N^{-1}T^{-1}\sigma_{SS}\right),\left\|\mathbf{R}^{1/2}\boldsymbol{\alpha}\right\|^{2}=o\left(\zeta_{1}^{-2}N^{-1}T^{-1}\sigma_{SS}^{2}\right)$
[…]
@@
Under Conditions (C1)-(C6), $\log N=o({T^{1/10}})$, and the alternative hypothesis (3), we have
@@ 5 Real data application
For the entire dataset ($T=165$), the $p$-values of the PY, MAX, COM, SS, SM, and CC tests are 7.70e-09, 0.0654, 1.54e-08, 9.48e-09, 0.0298, 3.60e-08, respectively. These results provide strong evidence that $\boldsymbol{\alpha}\neq\boldsymbol{0}$. Therefore, we employ a rolling window approach, applying tests to observations within each of the $165-T$ rolling windows. Each rolling window corresponds to a length-$T$ consecutive subsequence of $\{1,\ldots,165\}$, i.e., $\{\tau,\ldots,\tau+T-1\}$ for some $\tau\in\{1,\ldots,166-T\}$. Table 2 presents the rejection ratios of the tests at significance levels $\gamma=0.01,0.05$ and for three choices of $T=48,60,72$. These values of $T$ correspond to rolling windows covering 4, 5, or 6 years, respectively. Figure 3 show the $p$-value sequence of each test from 2005 to 2018 for $T=48,60,72$, respectively. Table 2 reveals that spatial-sign-based tests outperform corresponding least-squares-based tests, likely due to the non-normal distribution of residuals for most securities (see Figure S4 in Liu et al. (2023)). Additionally, sum-type test procedures exhibit better performance than max-type test procedures, which is unsurprising given the large number of nonzero alphas among the securities, as reported in Feng et al. (2022b). Finally, the newly proposed CC tests demonstrate nearly the best performance in most cases, highlighting the robustness of this method. These findings are consistent with theoretical results and simulation studies.
@@
Figure 3: $p$-value sequence of each test from 2005 to 2018 for US data.
(a) $T=48$
(b) $T=60$
(c) $T=72$
@@ Appendix A Proofs of all theorems
Under Conditions (C1)-(C4), we have $\max_{1\leq i\leq N}|\hat{d}_{i}^{2}-d_{i}^{2}|=O_{p}(\sqrt{\log N/T})$, where $\hat{{\bf D}}=\hbox{diag}\{\hat{d}_{1}^{2},\dots,\hat{d}_{N}^{2}\}$ and ${{\bf D}}=\hbox{diag}\{{d}_{1}^{2},\dots,{d}_{N}^{2}\}$.
@@
Under Conditions (C1)-(C5) and the null hypothesis, when $\min(T,N)\rightarrow\infty$, $T_{SS}\mathop{\rightarrow}\limits^{d}N(0,1)$.
@@ A.1 Proof of Theorem 1
$\quad\quad-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\hat{\boldsymbol{\theta}}-2\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$
$\quad\quad+2\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})^{-1/2}=\mathbf{0}.$
Let ${\boldsymbol{V}}_{t}={\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}$. Because ${\boldsymbol{f}}_{t}$ is independent of the errors $\varepsilon_{it}$, we have
$E(\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2})$
$=$ $E({\boldsymbol{f}}_{t}^{\top}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})$
$=$ $O(N)E({\boldsymbol{f}}_{t}^{\top}({\bf F}^{\top}{\bf F})^{-1}{\bf F}^{\top}{\bf I}_{T}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t})$
$=$ $O(NT^{-1}),$ (1)
where the last second equality holds because $\boldsymbol{\varepsilon}_{\cdot 1},\dots,\boldsymbol{\varepsilon}_{\cdot T}$ are independent and identically distributed random vectors, and $E(\boldsymbol{\varepsilon}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top})=E(\boldsymbol{\varepsilon}{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top})-E\{\boldsymbol{\varepsilon}({{\bf D}}^{-1}-\hat{{\bf D}}^{-1})\boldsymbol{\varepsilon}^{\top}\}=\{1+o_{p}(1)\}E(\varepsilon_{11}^{2}/d_{1}^{2}+\dots+\varepsilon_{1N}^{2}/d_{N}^{2}){\bf I}_{T}=O(N){\bf I}_{T}$. By the proof of Lemma A.3 in Feng et al. (2016) and $\|{\bf D}^{-1/2}\boldsymbol{\varepsilon}^{\top}\boldsymbol{V}_{t}\|^{2}=O_{p}(NT^{-1})$, we can obtain that $\|\hat{\boldsymbol{\theta}}\|=O_{p}(\zeta_{1}^{-1}T^{-1/2})$. According to the Taylor expansion, we can rewrite the above equation as
@@
$\quad\quad-\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}+\delta_{1t})=\mathbf{0},$
where $\delta_{1t}=O_{p}\big{\{}(-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}\|^{2}-\frac{1}{2}\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}\|^{2}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\hat{\boldsymbol{\theta}}+\hat{r}^{-1}_{t}\hat{{\bf U}}^{\top}_{t}\hat{{\bf D}}^{1/2}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t}-\hat{r}^{-2}_{t}\hat{\boldsymbol{\theta}}^{\top}\hat{{\bf D}}^{-1}\boldsymbol{\varepsilon}^{\top}{\boldsymbol{V}}_{t})^{2}\big{\}}$. Moreover, $\hat{r}^{-2}_{t}=\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}=\|{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}(1+\|(\hat{{\bf D}}^{-1/2}-{\bf D}^{-1/2}){\bf D}^{1/2}{\bf U}_{t}\|^{2}-2{\bf U}_{t}^{\top}(\hat{{\bf D}}^{-1/2}-{\bf D}^{-1/2}){\bf D}^{1/2}{\bf U}_{t})^{-1}={\color[rgb]{0,0,0}\|{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}_{\cdot t}\|^{-2}}\times\{1+o_{p}(1)\}=O_{p}(N^{-1})$. Hence, we have $\hat{r}^{-2}_{t}\|\hat{{\bf D}}^{-1/2}\boldsymbol{\varepsilon}^{\top}{\bf F}({\bf F}^{\top}{\bf F})^{-1}{\boldsymbol{f}}_{t}\|^{2}=O(T^{-1})$. Combined the above results and Condition (C3), we have $\delta_{1t}=O_{p}(T^{-1})$. Then, we have